# Visual-Friendly Concept Protection via Selective Adversarial Perturbations

### v2

@@ V-A Experimental Settings
TABLE I: Automatic and human evaluations across different methods, including their respective mean and standard deviation ($\pm$) on the VGGFace2 and CelebA-HQ datasets. All methods employed a perturbation size of $\epsilon$ = 8/255. In the HumanEval results, “Loss/Tie/Win” in “Visibility” indicates the voting rates for VCPro losses, ties, and wins in comparison to other methods regarding the stealthiness of their adversarial perturbations. “TI/DB” represents the percentage of images generated by Textual Inversion/DreamBooth identified as synthetic. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Methods Publication Visibility of Perturbations Textual Inversion DreamBooth HumanEval
FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$ Visibility (Loss/Tie/Win) TI $\uparrow$ DB $\uparrow$
VGGFace2 No Defense - - - - $3.61_{\pm 0.98}$ $0.24_{\pm 0.34}$ $3.95_{\pm 1.11}$ $0.38_{\pm 0.25}$ 64 : 25 : 11 - -
Mist arXiv23 $105.93_{\pm 9.76\hphantom{0}}$ $0.83_{\pm 0.03}$ $32.48_{\pm 1.12}$ $1.24_{\pm 0.41}$ $0.11_{\pm 0.29}$ $1.02_{\pm 0.11}$ $0.25_{\pm 0.28}$ 04 : 06 : 90 098.34 097.23
Anti-DreamBooth ICCV23 $\hphantom{0}96.24_{\pm 10.44}$ $0.82_{\pm 0.03}$ $32.43_{\pm 1.10}$ $1.18_{\pm 0.36}$ $0.05_{\pm 0.29}$ $1.03_{\pm 0.14}$ $0.26_{\pm 0.28}$ 03 : 09 : 88 097.38 100.00
PhotoGuard ICML23 $\hphantom{0}62.07_{\pm 8.15\hphantom{0}}$ $0.84_{\pm 0.03}$ $33.17_{\pm 1.21}$ $1.50_{\pm 0.54}$ $0.13_{\pm 0.34}$ $1.26_{\pm 0.30}$ $0.27_{\pm 0.31}$ 13 : 19 : 68 098.70 084.02
SDS(-) ICLR24 $\hphantom{0}47.48_{\pm 6.81\hphantom{0}}$ $0.81_{\pm 0.03}$ $32.91_{\pm 0.43}$ $2.09_{\pm 0.82}$ $0.07_{\pm 0.37}$ $2.74_{\pm 0.97}$ $0.33_{\pm 0.26}$ 13 : 16 : 71 094.56 77.13
MetaCloak CVPR24 $204.25_{\pm 54.52}$ $0.82_{\pm 0.02}$ $32.09_{\pm 0.63}$ $1.86_{\pm 0.62}$ $0.19_{\pm 0.33}$ $1.33_{\pm 0.41}$ $0.25_{\pm 0.32}$ 02 : 10 : 88 091.03 090.44
VCPro $\hphantom{0}27.04_{\pm 4.38\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.33_{\pm 1.79}$ $2.31_{\pm 0.78}$ $0.03_{\pm 0.32}$ $2.06_{\pm 0.70}$ $0.21_{\pm 0.34}$ 098.61 097.35
CelebA-HQ No Defense - - - $4.40_{\pm 0.86}$ $0.41_{\pm 0.22}$ $4.84_{\pm 0.40}$ $0.63_{\pm 0.13}$ 77 : 19 : 04 - -
Mist arXiv23 $\hphantom{0}78.32_{\pm 11.83}$ $0.86_{\pm 0.03}$ $33.75_{\pm 0.21}$ $1.45_{\pm 0.58}$ $0.26_{\pm 0.26}$ $1.04_{\pm 0.11}$ $0.44_{\pm 0.21}$ 05 : 04 : 91 096.29 99.99
Anti-DreamBooth ICCV23 $\hphantom{0}78.45_{\pm 11.66}$ $0.86_{\pm 0.03}$ $33.74_{\pm 0.21}$ $1.81_{\pm 1.20}$ $0.28_{\pm 0.25}$ $1.04_{\pm 0.15}$ $0.44_{\pm 0.21}$ 00 : 04 : 96 097.01 096.53
PhotoGuard ICML23 $\hphantom{0}45.18_{\pm 7.21\hphantom{0}}$ $0.88_{\pm 0.02}$ $35.08_{\pm 0.25}$ $1.81_{\pm 0.59}$ $0.30_{\pm 0.24}$ $1.23_{\pm 0.35}$ $0.48_{\pm 0.17}$ 02 : 22 : 76 094.25 094.31
SDS(-) ICLR24 $\hphantom{0}34.53_{\pm 6.16\hphantom{0}}$ $0.82_{\pm 0.03}$ $33.01_{\pm 0.28}$ $2.29_{\pm 0.70}$ $0.09_{\pm 0.28}$ $2.92_{\pm 0.74}$ $0.55_{\pm 0.15}$ 09 : 11 : 80 098.63 084.32
MetaCloak CVPR24 $161.93_{\pm 26.93}$ $0.86_{\pm 0.03}$ $33.12_{\pm 0.20}$ $1.47_{\pm 0.56}$ $0.25_{\pm 0.26}$ $1.59_{\pm 0.53}$ $0.44_{\pm 0.16}$ 00 : 04 : 96 096.02 095.48
VCPro $\hphantom{0}16.21_{\pm 2.89\hphantom{0}}$ $0.95_{\pm 0.01}$ $39.32_{\pm 0.40}$ $2.61_{\pm 1.00}$ $0.09_{\pm 0.28}$ $2.62_{\pm 0.95}$ $0.44_{\pm 0.19}$ 100.00 095.67
As image quality assessment remains a challenging problem, particularly with the lack of precise concept-specific automated metrics, we also provide human evaluation to assess the visual perception of protected data and the efficacy of safeguarding the target concepts.
Figure 3: Qualitative defense results for both the human face and non-face cases. For generating protected images, we use an $\epsilon$ value of 8/255 with Stable Diffusion v1-4. Each row displays the original images (No Defense), the protected images generated by PhotoGuard, Anti-DreamBooth, Mist, SDS(-), MetaCloak, and our method, along with their respective outcomes in Textual Inversion (TI) and DreamBooth (DB). Additionally, we include visualizations of adversarial perturbations represented as black-yellow images. For better detail, please zoom in.
V-B Main Results
Quantitative Results. Table I shows the automatic and human evaluation results for the visibility of protected images and the final protection effectiveness of different methods.
For visibility of protected images, we evaluate the average FID/SSIM/PSNR differences between all protected images and their corresponding original images. The statistics show that the protected images generated by our method are more faithful to the original images, and the least variance of FID shows that we also have a stable effect. Compared with our baseline Anti-DreamBooth, we can max reduce 69.20 in FID, up to 0.09 in SSIM, and up to 5.58 in PSNR. Compared with current SOTA methods PhotoGuard and SDS (-), our image quality is still significantly improved. Although SDS (-) achieves a low FID, it introduces noticeable brightness increases in the perturbations, leading to poor SSIM performance.
Protected images against unauthorized AI generation by improving training costs and reducing final AI-generated image quality. For final protection effectiveness, we still have significant relative decreases in full-image quality metric LIQE compared to the no defense (from “good/perfect” to “poor”), again achieving a protective effect, which is similarly confirmed by the manual evaluation later on human evaluation and qualitative results. Considering that the full-image quality metric is not fairly for us, we also add a face region quality metric CLIP-FACE. In CLIP-FACE, we achieve the best protection in most cases. Both metrics demonstrate that we have reduced visibility against perturbations while still having a strong protection effect.
Qualititative Results. Fig. 3 illustrates protected images and protective effects on individual identity information and landscapes of different methods under Textual Inversion and DreamBooth. We also visualize adversarial perturbations by regularizing all perturbations to the range [0,1] and then visualizing using a specific colormap.
Under the protection of our method, the personalization methods fail to generate visual-fidelity results. As shown in the case of important information such as faces, landmark landscapes, or buildings, our adversarial perturbation is less noticeable, especially in the background, and effectively distorts textures and features in crucial areas. This prevents the unauthorized use of personal identification information or copyrighted material, making the perturbation more subtle while still serving its protective purpose. This approach strikes a better balance between safeguarding the content and maintaining its quality, thereby reducing user concerns about image degradation after protection. PhotoGuard, Mist, Anti-DreamBooth, and MetaCloak tend to add strange textures to the whole image and show obvious adversarial perturbations (the second, third, fourth, and sixth columns). SDS(-) can cause generated images to be too blurry and has limited protection for the DreamBooth. Its protected image is like adding strange circular blobs to the original image (the fifth column). Visualization of its perturbations also supports the idea that excessive brightness leads to lower SSIM values. These approaches serve as a protective measure, ensuring that certain image regions exhibit unique textures. The presence of such textures can strongly suggest infringement. We also find that the protection difficulty of Textual Inversion is far less than that of DreamBooth because DreamBooth usually fine-tunes most parameters in Stable Diffusion. At the same time, Textual Inversion only adds a few new word embeddings. It is also supported in Table I.
User Study. We evaluate the performance of VCPro in the concept protection task with five selected methods, i.e., PhotoGuard [12], Mist [14], Anti-DreamBooth [8], SDS(-)[11], and MetaCloak [10]. A total of 50 participants took part in the survey, and they all use social media proficiently. Participants comprise 58% male and 42% female, with ages ranging from 18 to 55 years (mean: 24.5). All participants have normal color vision and social media proficiency relevant to our image protection context. Educational backgrounds span from undergraduate to doctoral level (undergraduate: 54% graduate: 26%, doctoral: 4%). Professional expertise was distributed across computer science (50%), electronic information technology (36%), and visual arts (14%). For quality control, all participants completed preliminary training and visual acuity testing. We implemented consistency checks through randomized question repetition with shuffled options, requiring participants to maintain at least 95% response consistency for inclusion in the final analysis.
We randomly sampled protected images and generated images from all 100 identities in VGGFace2 and CelebA-HQ and six groups of non-face data. The study consists of three parts, including visibility of adversarial perturbations (2064 valid votes), protection effects under Textual Inversion (1996 valid votes), and protection effects under DreamBooth (1990 valid votes).
User Study I. In the survey on the visibility of adversarial perturbations, participants were shown a VCPro-protected image alongside a protected image generated by another method based on the same original image. They were then asked to vote on the visual quality, choosing between “A wins”, “tie”, or “B wins”. The visual quality of VCPro was significantly better than that of other methods. As shown in Table I, our VCPro significantly outperformed all baseline methods in direct comparisons. VCPro received between 68% and 96% of participant votes when compared against competing methods. To assess the reliability of our user study results, we calculated the Kendall coefficient [41] as 0.71 ($p<0.05$), indicating substantial inter-rater agreement among participants.
User Study II. In the survey on protection efficacy under Textual Inversion, participants were given an original image $x$ and asked to determine whether the image generated through Textual Inversion, trained on protected image $x^{\prime}$ from different methods, was synthetic. As shown in Table I, all protection methods received at least 91% of synthetic votes, with VCPro demonstrating the most effective protection as acknowledged by the majority of participants. To evaluate the consistency of the user study, we calculated Cohen’s Kappa [42], which yielded a value of 0.88 ($p<0.05$), indicating a strong agreement among participants.
@@ V-B Main Results
User Study III. In the survey on protection efficacy under DreamBooth, as shown in Table 1, participants recognized an average 97.01% protective success rate of VCPro. To measure the consistency of the study, we calculated Cohen’s Kappa, which resulted in a value of 0.88 ($p<0.05$).
TABLE II: Automatic evaluation results on ablation experiments with the corresponding mean and standard deviation ($\pm$). All methods use $\epsilon$ = 8/255 on VGGFace2 dataset. ↑ means the higher the better, and vice versa. The best result in each column is in bold. Here Input-Mask is to constrain the adversarial perturbations directly on the input image to be only within the mask, we are constraining on the optimization objective, and the adversarial perturbations will be present in the whole image.
Methods Visibility of Perturbations Textual Inversion DreamBooth
No Defense - - - $3.61_{\pm 0.98}$ $0.24_{\pm 0.34}$ $3.95_{\pm 1.11}$ $0.38_{\pm 0.25}$
Anti-DreamBooth $96.24_{\pm 10.44}$ $0.82_{\pm 0.03}$ $32.43_{\pm 1.10}$ $1.18_{\pm 0.36}$ $0.05_{\pm 0.29}$ $1.03_{\pm 0.14}$ $0.26_{\pm 0.28}$
Anti-DreamBooth+RAL $68.18_{\pm 8.58\hphantom{0}}$ $0.83_{\pm 0.03}$ $32.89_{\pm 1.25}$ $1.62_{\pm 0.57}$ $0.03_{\pm 0.31}$ $1.53_{\pm 0.53}$ $0.18_{\pm 0.31}$
Anti-DreamBooth+RAL (Small Mask) $58.44_{\pm 9.34\hphantom{0}}$ $0.84_{\pm 0.03}$ $33.09_{\pm 1.27}$ $2.11_{\pm 0.82}$ $0.05_{\pm 0.33}$ $2.21_{\pm 0.66}$ $0.27_{\pm 0.32}$
Anti-DreamBooth+LMLS $30.08_{\pm 5.22\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.03_{\pm 1.83}$ $2.52_{\pm 0.82}$ $0.17_{\pm 0.34}$ $2.16_{\pm 0.65}$ $0.30_{\pm 0.30}$
VCPro (w/o Pull) $17.21_{\pm 3.13\hphantom{0}}$ $0.94_{\pm 0.03}$ $36.79_{\pm 2.34}$ $2.74_{\pm 0.82}$ $0.14_{\pm 0.33}$ $2.55_{\pm 0.92}$ $0.35_{\pm 0.32}$
VCPro $27.04_{\pm 4.38\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.33_{\pm 1.79}$ $2.31_{\pm 0.78}$ $0.03_{\pm 0.32}$ $2.06_{\pm 0.70}$ $0.21_{\pm 0.34}$
Anti-DreamBooth+Input-Mask $41.27_{\pm 8.76\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.15_{\pm 2.02}$ $2.05_{\pm 0.76}$ $0.10_{\pm 0.32}$ $2.03_{\pm 0.80}$ $0.23_{\pm 0.33}$
VCPro+Input-Mask $18.61_{\pm 3.19\hphantom{0}}$ $0.93_{\pm 0.03}$ $36.61_{\pm 2.25}$ $2.41_{\pm 0.78}$ $0.07_{\pm 0.34}$ $2.68_{\pm 1.05}$ $0.31_{\pm 0.32}$
[…]
V-C Ablation Study
As shown in Fig. 4 and Table II, we conducted ablation experiments of two modules in VCPro: Regional Adversarial Loss (RAL) in Sec. IV-B and Lagrangian Multiplier-based Loose Solution (LMLS) in Sec. IV-C. Anti-DreamBooth is our baseline.
Moreover, only push-loss without pull-loss in RAL (VCPro (w/o Pull) in Fig. 4 and Table II) will reduce the training loss gradient feedback about the protection effect, while the loss that constrains the visibility of the perturbations is not affected resulting in decreased visibility of the perturbations and the protection effect.
V-D Perturbation Size Influence
Perturbation size $\epsilon$ controls the maximum allowable change in pixel values of adversarial perturbation. As shown in Fig. 5 and Table III, different levels of adversarial perturbation size noticeably influence the protection outcomes: Compared to low $\epsilon$, large $\epsilon$ presents worse invisibility of adversarial perturbations while the more obvious protection effects.
TABLE III: The influence of different perturbation size $\epsilon$ on VGGFace2 datasets. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@ V-D Perturbation Size Influence
V-E Hyper-parameter Analysis

### v3

@@ Introduction
No Defense 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
Anti-DB 96.2±10.4 0.82±0.03 32.4±1.1 1.18±0.36 0.05±0.29 1.03±0.14 0.26±0.28
Anti-DB+RAL 68.2±8.6 0.83±0.03 32.9±1.3 1.62±0.57 0.03±0.31 1.53±0.53 0.18±0.31
Anti-DB+RAL (Small Mask) 58.4±9.3 0.84±0.03 33.1±1.3 2.11±0.82 0.05±0.33 2.21±0.66 0.27±0.32
Anti-DB+LMLS 30.1±5.2 0.90±0.03 35.0±1.8 2.52±0.82 0.17±0.34 2.16±0.65 0.30±0.30
VCPro (w/o Pull) 17.2±3.1 0.94±0.03 36.8±2.3 2.74±0.82 0.14±0.33 2.55±0.92 0.35±0.32
VCPro 27.0±4.4 0.90±0.03 35.3±1.8 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
Anti-DB+Input-Mask 41.3±8.8 0.90±0.03 35.2±2.0 2.05±0.76 0.10±0.32 2.03±0.80 0.23±0.33
VCPro+Input-Mask 18.6±3.2 0.93±0.03 36.6±2.3 2.41±0.78 0.07±0.34 2.68±1.05 0.31±0.32
Table 2: Ablation study on VGGFace2 with $\epsilon=8/255$. Input-Mask constrains perturbations within the mask region on input images (vs. our method which constrains the optimization objective, allowing perturbations across the entire image). $\uparrow$$\downarrow$ indicates higher/lower is better. Bold: best performance.
Main Results
Quantitative Results. Table 1 shows automatic and human evaluation results for protected image visibility and protection effectiveness. For visibility metrics, our method generates protected images most faithful to originals with stable performance (lowest FID variance). Compared to Anti-DreamBooth baseline, we achieve up to 69.20 FID reduction, 0.09 SSIM improvement, and 5.58 PSNR increase. Our method significantly outperforms current SOTA methods PhotoGuard and SDS (-) in image quality, though SDS (-) achieves low FID by introducing noticeable brightness increases that hurt SSIM performance.
For protection effectiveness, we achieve significant quality degradation in full-image metric LIQE (from “good/perfect” to “poor”) and excel in face region quality metric CLIP-FACE. Both metrics demonstrate reduced visibility against perturbations while maintaining strong protection.
Qualitative Results. Fig. 3 shows protected images and effects on identity information and landscapes under Textual Inversion and DreamBooth. Adversarial perturbations are visualized using [0,1] normalization with colormaps. PhotoGuard, Mist, Anti-DreamBooth, and MetaCloak add obvious strange textures throughout images. SDS(-) causes excessive blurriness with limited DreamBooth protection, adding circular blob artifacts. Our approach achieves subtler perturbations while ensuring protective effectiveness. Our method prevents personalization methods from generating high-fidelity results. For faces, landmarks, and buildings, our perturbations are less noticeable (especially in backgrounds) while effectively distorting crucial textures and features. This prevents unauthorized use of personal or copyrighted material while maintaining better quality balance.
Textual Inversion protection is easier than DreamBooth since DreamBooth fine-tunes most Stable Diffusion parameters while Textual Inversion only adds word embeddings, as supported by Table 1.
User Study. We evaluate VCPro against five methods: PhotoGuard (salman2023raising), Mist (liang2023mist), Anti-DreamBooth (le_etal2023antidreambooth), SDS(-) (xue2024toward), and MetaCloak (liu2024metacloak). 50 participants (58% male, 42% female, ages 18-55, mean: 24.5) with social media proficiency participated. We randomly sampled protected and generated images from 100 identities in VGGFace2 and CelebA-HQ plus six non-face groups. The study includes three parts: perturbation visibility (2064 valid votes), Textual Inversion protection (1996 valid votes), and DreamBooth protection (1990 valid votes).
User Study I. Participants compared VCPro-protected images with other methods’ results, voting on visual quality (“A wins”, “tie”, or “B wins”). VCPro received 68%-96% of votes against competing methods with Kendall coefficient of 0.71 ($p<0.05$), indicating substantial inter-rater agreement.
User Study II. Participants determined whether images generated via Textual Inversion from protected images were synthetic. All methods achieved $\geq$ 91% synthetic recognition, with VCPro showing most effective protection. Cohen’s Kappa = 0.88 ($p<0.05$) indicates strong participant agreement.
@@
User Study III. For DreamBooth protection evaluation, participants recognized 97.01% protective success rate for VCPro. Cohen’s Kappa = 0.88 ($p<0.05$) confirms study consistency.
@@
Table 3: The influence of different perturbation size $\epsilon$ on VGGFace2 datasets. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Ablation Study
As shown in Table 3, we conduct ablation experiments on two VCPro modules: Regional Adversarial Loss (RAL) and Lagrangian Multiplier-based Loose Solution (LMLS), using Anti-DreamBooth as baseline. Anti-DreamBooth+RAL reduces perturbation visibility with protection concentrated in mask areas, resulting in slight LIQE decrease but stable CLIP-FACE performance. Smaller masks further reduce perturbation visibility but limit protection range. Anti-DreamBooth+LMLS reduces noise visibility and preserves high-frequency components, but final protection creates streaky textures across the entire image. This confirms that adversarial perturbations in specific spatial/frequency domains directly affect corresponding protection domains. VCPro without pull-loss (VCPro (w/o Pull)) reduces training gradient feedback for protection while maintaining perturbation visibility constraints, decreasing both visibility and protection effectiveness. Comparing direct mask-constrained updates (Input-Mask) with our loss-guided approach shows both achieve regional protection. Direct masking improves pixel-level metrics (SSIM/PSNR) but makes perturbations more obvious with high FID. Our combined approach achieves better visual perception while maintaining target protection.
Perturbation Size Influence.
Perturbation size $\epsilon$ controls the maximum allowable change in pixel values of adversarial perturbation. As shown in Table 4, different levels of adversarial perturbation size noticeably influence the protection outcomes: Compared to low $\epsilon$, large $\epsilon$ presents worse invisibility of adversarial perturbations while the more obvious protection effects. For Textual Inversion, when the perturbation size is low, the image preserves the facial area with alterations in facial texture and feature distribution. Upon reaching a perturbation size of 16/255, the facial areas experience complete degradation. The results under DreamBooth and quantitative experiment show a similar trend. Compared with DreamBooth, Textual Inversion is easier to achieve concept protection.
Qualitative ablation experiments, hyper-parameter (training iterations, $c$, $\alpha$), analysis of adversary settings, and frequency domain analysis please see supplementary materials.
Conclusion
Future Works and Limitations. Our future efforts will focus on finding efficient methods to produce protected images swiftly while maintaining high visual quality. By doing so, we aim to significantly enhance the overall user experience and ensure that our solutions meet the highest standards of both functionality and aesthetics.
Ethical Considerations. VCPro empowers individuals against AI content power imbalances. Technical measures complement limited legal protections. Dual-use potential—adversaries could theoretically exploit VCPro to evade accountability. However, the primary threat today is unauthorized personalization with limited user consent, and empowering individuals is essential given current power asymmetries.
Acknowledgments
This work was partly supported by the National Natural Science Foundation of China under No. 62572458, and the Innovation Funding of ICT, CAS under Grant No. E561160.
Appendix
Visual-Friendly Concept Protection via Selective Adversarial Perturbations
Appendix A Detailed Experimental Settings
Datasets. We used the CelebA-HQ (karras2017progressive) and VGGFace2 (cao2018vggface2) datasets, following Anti-DreamBooth (le_etal2023antidreambooth) and MetaCloak (liu2024metacloak), to validate the effectiveness of our method in privacy protection. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of $30,000$ celebrity face images. VGGFace2 is a large-scale dataset with over $3.3$ million face images from $9,131$ unique identities. We randomly selected 50 identities from each dataset, ensuring that each had at least 15 images with a resolution exceeding 500 × 500 pixels. We also provide other non-face cases (artwork and landscape) in the paper.
Training details. The availability of Stable Diffusion pre-trained weights on Hugging Face  [^1] https://huggingface.co/ has significantly advanced research within the community. Here, our experiments primarily focus on the Stable Diffusion v1-4 version. For image pre-processing, we center crop and resize images to a resolution $512\times 512$. During protected image generation, we use $720$ iterations, step size $\alpha$ is $1/255$, and adversarial perturbation size $\epsilon$ is $8/255$. We test the protection performance on two types of SOTA personalization methods, i.e., Textual Inversion (TI) (gal2022image), and DreamBooth (DB) (ruiz2023dreambooth). During the training of the Textual Inversion, the constant learning rate is set to $5\times 10^{-4}$ with $3000$ optimization steps and batch size $1$. In DreamBooth, the constant learning rate is set to $5\times 10^{-7}$ with $1000$ optimization steps and batch size is $2$. We use a guidance scale of $7.5$ and $50$ denoising steps at the generation phase with a testing prompt. We set $c=0.1$ and $\alpha=0.5$ in practice. All experiments are complete on NVIDIA A100 GPU 40GB.
Baselines. We compare our method with five SOTA baselines, i.e., Mist (liang2023adversarial), Anti-DreamBooth (referred to as Anti-DB) (le_etal2023antidreambooth), PhotoGuard (salman2023raising), SDS(-) (xue2024toward) and MetaCloak (liu2024metacloak). In implementations, we set $\epsilon=8/255$ as the $\ell_{\infty}$ perturbation size, $\alpha=1/255$ as the step size for fair comparison.
  • Mist aims to maximize the training loss of Latent Diffusion Models (LDMs) while minimizing a textual loss between the protected image and a dummy image in the image encoder layer during protected image generation. In our implementation, following the original paper (liang2023mist), we set the loss weight for the textual loss to $1\times 10^{-4}$ and run 720 iterations during protected image generation.
  • PhotoGuard targets diffusion-based inpainting/editing methods and offers two variants: one that attacks the VAE-encoder of LDMs, and another that attacks the final image outputs. We select the VAE-encoder variant due to its superior visual quality and set the training iterations to 200, as recommended in the original paper (salman2023raising).
  • SDS(-) minimizes the SDS loss (poole2022dreamfusion) to optimize adversarial data, achieving the best visual quality among all versions presented in its original paper (xue2024toward). We set the training iterations to 100, following (xue2024toward).
  • MetaCloak utilizes meta-learning to enhance the robustness of adversarial data against various transformations. We set the number of surrogate models to 5, the unrolling number to 1, the sample batch size to 1, and the training iterations to 4000, based on the recommendations in (liu2024metacloak).
Metrics. To measure the visual perception of protected data, we calculate the FID (heusel2017gans), SSIM (wang2004image), and PSNR between protected and original images. The goal of the protection is to reduce the visual quality of generated images when adversaries use Textual Inversion and DreamBooth for personalization generation. So we use a full image quality evaluation metric LIQE (zhang2023blind) and propose a face regional quality evaluation metric CLIP-FACE (liu2024metacloak) as automatic evaluation metrics to assess protection effectiveness. LIQE is an advanced blind image quality assessment tool that measures image quality on a five-point scale: $c\in C=\{1,2,3,4,5\}=\{\text{``bad''},\text{``poor''},\text{``fair''},\text{``good''},\text{``perfect''}\}$. CLIP-FACE is based on CLIP-IQA (wang2023exploring) for visual quality by considering additional class information. Specifically, we calculate the CLIP score difference between “good face” and “bad face”. Considering the large number of images generated by Textual Inversion or DreamBooth, fine-grained labeling of facial regions in every image and then evaluating them is expensive. CLIP-FACE uses the multimodal understanding capability of CLIP and can focus on the facial regions in a coarse-grained manner. To assess the efficacy of the protection mechanism, we generate $16$ images for each trained Textual Inversion model and DreamBooth model with testing prompt “a photo of skssks person” or “a photo of skssks object”. As image quality assessment remains a challenging problem, particularly with the lack of precise concept-specific automated metrics, we also provide human evaluation to assess the visual perception of protected data and the efficacy of safeguarding the target concepts.
Appendix B Ablation Study
Method Visibility TI DB
[…]
No Defense 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
Anti-DB 96.2±10.4 0.82±0.03 32.4±1.1 1.18±0.36 0.05±0.29 1.03±0.14 0.26±0.28
Anti-DB+RAL 68.2±8.6 0.83±0.03 32.9±1.3 1.62±0.57 0.03±0.31 1.53±0.53 0.18±0.31
Anti-DB+RAL (Small Mask) 58.4±9.3 0.84±0.03 33.1±1.3 2.11±0.82 0.05±0.33 2.21±0.66 0.27±0.32
Anti-DB+LMLS 30.1±5.2 0.90±0.03 35.0±1.8 2.52±0.82 0.17±0.34 2.16±0.65 0.30±0.30
VCPro (w/o Pull) 17.2±3.1 0.94±0.03 36.8±2.3 2.74±0.82 0.14±0.33 2.55±0.92 0.35±0.32
VCPro 27.0±4.4 0.90±0.03 35.3±1.8 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
Anti-DB+Input-Mask 41.3±8.8 0.90±0.03 35.2±2.0 2.05±0.76 0.10±0.32 2.03±0.80 0.23±0.33
VCPro+Input-Mask 18.6±3.2 0.93±0.03 36.6±2.3 2.41±0.78 0.07±0.34 2.68±1.05 0.31±0.32
Table 3: Ablation study on VGGFace2 with $\epsilon=8/255$. Input-Mask constrains perturbations within the mask region on input images (vs. our method which constrains the optimization objective, allowing perturbations across the entire image). $\uparrow$$\downarrow$ indicates higher/lower is better. Bold: best performance.
[…]
As shown in Fig. 4 and Table 3, we conducted ablation experiments of two modules in VCPro: Regional Adversarial Loss (RAL) and Lagrangian Multiplier-based Loose Solution (LMLS). Anti-DreamBooth is our baseline.
Moreover, only push-loss without pull-loss in RAL (VCPro (w/o Pull) in Fig. 4 and Table 3) will reduce the training loss gradient feedback about the protection effect, while the loss that constrains the visibility of the perturbations is not affected resulting in decreased visibility of the perturbations and the protection effect.
Appendix C Perturbation Size Influence
Perturbation size $\epsilon$ controls the maximum allowable change in pixel values of adversarial perturbation. As shown in Fig. 5 and Table 4, different levels of adversarial perturbation size noticeably influence the protection outcomes: Compared to low $\epsilon$, large $\epsilon$ presents worse invisibility of adversarial perturbations while the more obvious protection effects.
Budget Visibility of Perturbations Textual Inversion DreamBooth
0 - - - 3.61±0.98 00.24±0.34 3.95±1.11 0.38±0.25
Table 4: The influence of different perturbation size $\epsilon$ on VGGFace2 datasets. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Appendix D Hyper-parameter Analysis